TheSmartOne (thesmartone):

Find value(s) of parameter c (if exists) that make f continuous everywhere.

TheSmartOne (thesmartone):

\(\Large f(x) = \frac{2}{x^2+4x+c}\)

TheSmartOne (thesmartone):

There was a question before it with the same instructions \(\Large f(x) = \frac{2}{x+c}\) I feel like the answers to both of them are there is no c value that will make f continuous everwhere.

OpenStudy (unklerhaukus):

Is there some restriction on the domain of f ?

TheSmartOne (thesmartone):

No

OpenStudy (unklerhaukus):

what are the solution to x^2+4x+c = 0 ? when are these solutions not real numbers?

TheSmartOne (thesmartone):

x^2 + 4x + 4 = 0 (x + 2)^2 = 0 x = -2 c = (-2)^2 = 4 so whenever c > 4 the solutions to it will not be real numbers

OpenStudy (unklerhaukus):

when c>4 the denominator cannot be zero

TheSmartOne (thesmartone):

I made c = 4 (because perfect square trinomials) and then calculated x and then found the value of c again and I'm going crazy x'D But when c > 4 the solutions will not be real numbers

TheSmartOne (thesmartone):

Ohhhh o: That makes a lot of sense! Thanks Unkle! :D